OpenStudy (anonymous):

Find the value by referring to the graph of the trigonometric function. If the value is undefined, write undefined. tan(-3/2 pi)

OpenStudy (jdoe0001):

check your Unit Circle for the angle \(\bf \cfrac{3\pi}{2}\) keeping in mind that \(\bf tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\implies \cfrac{y}{x}\)

OpenStudy (jdoe0001):

\(\bbf -\cfrac{3\pi}{2}\ rather, missed the negative part, keep in mind that negative angles are the ones that go CLOCKWISE 

OpenStudy (jdoe0001):

\(\bf -\cfrac{3\pi}{2}\) rather